Represent each of these relations on {1, 2, 3} with a matrix (with the elements


of this set listed in increasing order).


a) {(1, 1), (1, 2), (1, 3)}


b) {(1, 2), (2, 1), (2, 2), (3, 3)}


c) {(1, 1), (1, 2), (1, 3), (2, 2), (2, 3), (3, 3)}


d) {(1, 3), (3, 1)}

1
Expert's answer
2021-12-02T18:39:24-0500

Let us represent each of these relations on "\{1, 2, 3\}" with a matrix (with the elements of this set listed in increasing order).


a) For the relation "\{(1, 1), (1, 2), (1, 3)\}" the matrix is the following:


"\begin{pmatrix}
1 & 1& 1\\
0 & 0 & 0\\
0 & 0 & 0
\end{pmatrix}"


b) "\{(1, 2), (2, 1), (2, 2), (3, 3)\}"


"\begin{pmatrix}
0 & 1 & 0\\
1 & 1 & 0\\
0 & 0 & 1
\end{pmatrix}"


c) "\{(1, 1), (1, 2), (1, 3), (2, 2), (2, 3), (3, 3)\}"


"\begin{pmatrix}
1 & 1& 1\\
0 & 1 & 1\\
0 & 0 & 1
\end{pmatrix}"


d) "\{(1, 3), (3, 1)\}"


"\begin{pmatrix}
0 & 0 & 1\\
0 & 0 & 0\\
1 & 0 & 0
\end{pmatrix}"